b) Your company is considering adding a new product line. The new equipment cost is $360,000 per year with a salvage value of $40,000 at the end of year 5. The variable cost per unit of the new product line will be $14.55 and the overhead cost per year to be $48,000. If the company uses a 5-year planning horizon and the price of the product is $49.99, how many units are needed to break even? No interest = 0
Answer #1

In the case of break even, total cost = total revenue.

Let us assume x units are required to be produce to break even.

Let us calculate the total costs in 5 years. Cost of equipment = purchase value - salvage value = $360,000 - $40,000 = $320,000.

Total variable cost = variable costs per unit * no. of units = $14.55 * x = $14.55x

Total overhead costs in 5 years = $48,000 * 5 = $240,000.

Total revenue = $49.99x

So, at the break even point:

Total cost = Total revenue

or, 320,000 + 14.55x + 240,000 = 49.99x

or, 560,000 = 49.99x - 14.55x

or, 560,000 = 49.99x - 14.55x = 35.44x

or, x = 560,000/35.44 = 15801 units.